On a 200 km bike cycle ride, you ride the first 90 km at an average speed of 20 m/s. What must your average speed over the next 40 km be to have your average speed for the total 200 km be?

 (a)  60 m/s?

 (b) Given this average speed for the first 90 km, can you possibly attain an average speed of

    80 m/s for the total 200 km ride? Explain.

 

Please reply as fast as possible


1
Expert's answer
2021-01-28T17:09:23-0500

a)


"v=\frac{s}{t}\\60=\frac{130000}{\frac{90000}{20}+\frac{40000}{u}}\\u=-17\frac{m}{s}"

b) No.


"v=\frac{s}{t}\\80=\frac{130000}{\frac{90000}{20}+\frac{40000}{u}}\\u=-14\frac{m}{s}"